OpenStudy (anonymous):

Write the equation of the line perpendicular to 9x + 2y = 18 and passes through the point (1, 4)

OpenStudy (anonymous):

PLEASE HELP

OpenStudy (saifoo.khan):

\[\Large \begin{array}l\color{red}{\text{C}}\color{orange}{\text{H}}\color{#9c9a2e}{\text{I}}\color{green}{\text{L}}\color{blue}{\text{L}}\color{purple}{\text{ }}\color{purple}{\text{D}}\color{red}{\text{U}}\color{orange}{\text{D}}\color{#9c9a2e}{\text{E}}\color{green}{\text{!}}\color{blue}{\text{}}\end{array}\]

OpenStudy (anonymous):

if you dont want to help just say so

OpenStudy (saifoo.khan):

lol, why wont i?

OpenStudy (saifoo.khan):

i've doing that all this time. :p

OpenStudy (anonymous):

whyd you tell me to chill?

OpenStudy (saifoo.khan):

lol, thts to calm down!

OpenStudy (anonymous):

probably because you used caps to ask for help

OpenStudy (saifoo.khan):

Alan are u same from the movie Hangover?

OpenStudy (anonymous):

use your inside internet voice.

OpenStudy (saifoo.khan):

hmm, right! ^ ^

OpenStudy (anonymous):

thats the only way ppl will help me on here

OpenStudy (anonymous):

can you help me or not??

OpenStudy (saifoo.khan):

Lol, what if we dont?? then?

OpenStudy (anonymous):

then ill try yo find someone else..but I would love for you to help me

OpenStudy (anonymous):

sooo you need negative reciprecals. Hold on I forgot how to do this

OpenStudy (anonymous):

ok..

OpenStudy (saifoo.khan):

LOl, If u want me to help, i will surely! :)

OpenStudy (anonymous):

saifoo. Help lol. I wanna know how to do this too!

OpenStudy (saifoo.khan):

alright, then i will do this step by stel.

OpenStudy (saifoo.khan):

step*

OpenStudy (saifoo.khan):

the slope of a perp line follow this formula: \[\Large m_1 * m_2 = -1\] now insert the values.. the slope for the real line is: -9/2 so. \[\frac{-9}{2} *m_2 = -1\] \[m_2 = \frac92\] so here we have the slope for that.

OpenStudy (anonymous):

the Y intercept is the same right?

OpenStudy (saifoo.khan):

now, insert this in y=mx +b \[4 = \frac{9*1}{2} + b \]\[b = \frac{-1}{2}\] so now u have the y intercpt too.

OpenStudy (anonymous):

I guess I forgot how to use y=mx + b. I'm sensing with my powers that the 4 is from his given point, and so the 1 too must be from his given point. So you're only adding the 9/2 and + b?

OpenStudy (saifoo.khan):

now, just insert the slope & y-intercpt: \[y = \frac{-9x}{2} -\frac{1}{2}\] finally take the lcm and the cross multiply.. finally your equation will look like: \[\Huge 2y = -9x - 1\]

OpenStudy (anonymous):

what. How did you get -1/2 for b >.>

OpenStudy (anonymous):

the answers are either: a. 9x-2y=1 b.9x+2y=17 c.2x-9y=-34 d.2x+9y=38

OpenStudy (saifoo.khan):

i inserted the slope and the new cordinates.. to find the y-intercpt.

OpenStudy (saifoo.khan):

A for apple is the answer.

OpenStudy (anonymous):

okee doke.. thanx!

OpenStudy (saifoo.khan):

welkx. i never did such LONG working in my 2mnths here!

OpenStudy (anonymous):

wait you inserted 9/2... and the coordinates 4 for y and 1 for x lol so how did you get -1/2

OpenStudy (saifoo.khan):

lol, i did the working there.. u can see that.

OpenStudy (anonymous):

I can see you multiplying through by 2 to get 8 on the left and 9 on the right, subtract to get b = 1

OpenStudy (saifoo.khan):

huh?

OpenStudy (anonymous):

OH

OpenStudy (anonymous):

4-4.5 = -1/2. this is why I fail at math

OpenStudy (anonymous):

SAIFOO OFF THE TOP OF YOUR HEAD WHAT IS THE FOUR STEP METHOD TO INVERSE A FUNCTION????

OpenStudy (saifoo.khan):

LOL.

OpenStudy (anonymous):

:3 tell meNAOW

OpenStudy (saifoo.khan):

sorry, where u got stuck?

OpenStudy (anonymous):

>.> I'm working on a different question. Don't worry I got the answer right here.

OpenStudy (saifoo.khan):

O cool. the last step i do is chng the y to x.

OpenStudy (anonymous):

isn't that the first step... you change back to y at the end